# Tensor networks enable the calculation of turbulence probability distributions

### v1

@@ Abstract
Predicting the dynamics of turbulent fluid flows has long been a central goal of science and engineering. Yet, even with modern computing technology, accurate simulation of all but the simplest turbulent flow-fields remains impossible: the fields are too chaotic and multi-scaled to directly store them in memory and perform time-evolution. An alternative is to treat turbulence probabilistically, viewing flow properties as random variables distributed according to joint probability density functions (PDFs). Turbulence PDFs are neither chaotic nor multi-scale, but are still challenging to simulate due to their high dimensionality. Here we show how to overcome the dimensionality problem by parameterising turbulence PDFs into an extremely compressed format known as a “tensor network” (TN). The TN paradigm enables simulations on single CPU cores that would otherwise be impractical even with supercomputers: for a $5+1$ dimensional PDF of a chemically reactive turbulent flow, we achieve reductions in memory and computational costs by factors of $\mathcal{O}(10^{6})$ and $\mathcal{O}(10^{3})$, respectively, compared to standard finite difference algorithms. A future path is opened towards something heretofore regarded as infeasible: directly simulating high-dimensional PDFs of both turbulent flows and other chaotic systems that are useful to describe probabilistically.
@@ Introduction
Despite the simple and deterministic physical laws governing it, turbulence remains an inherently complex and chaotic phenomenon. It is characterised by large numbers of eddies interacting in intricate and nonlinear ways across wide ranges of spatial and temporal scales, leading to the emergence of chaos. Making matters worse, practically important turbulent flows (e.g. fuel-oxidiser mixtures in combustion) often involve multiple chemically-reacting species, which introduces additional nonlinearities and scales. The presence of chaos prohibits predicting the exact dynamics of turbulent flow-fields over long periods of time, while the multi-scaled nature of the flow-fields makes their simulation immensely expensive, due to the need to solve sets of coupled partial differential equations (PDEs) on very fine grids.
However, for practical applications it is rarely necessary to know the precise state of a turbulent flow-field at every point in space-time. Rather, one is typically more interested in far slower-varying statistical quantities where the fluctuations are averaged out (such as the lift and drag of an aeroplane, or the rate of product formation in a chemical process). In the statistical description of turbulence, variables like velocities $\mathbf{U}$, chemical mass-fractions $\Phi_{\alpha}$, temperatures, etc., are treated as random variables (RVs) distributed according to some one-point, one-time joint probability density function (PDF) [1]
@@
across space $\mathbf{x}$ and time $t$, with $\mathbf{u},\varphi_{1}$ being sample-space variables corresponding to $\mathbf{U},\Phi_{1}$. The trajectory of $f$ completely describes the one-point, one-time statistics of the flow dynamics [2]; which are the central quantities of interest in practical engineering calculations.
@@
The time-evolution of $f$ is modelled by Fokker-Planck PDEs that are straightforward to derive [3, 4, 5], but hard to solve. If $f$ is $d$-dimensional, assigning $M$ points for each dimension results in a total of $M^{d}$ gridpoints. Given that $d$ can be as high as $\mathcal{O}(10^{3})$ in realistic flows [6, 7], direct schemes like finite differences (FD) or volumes were long ago dismissed as computationally infeasible [8] due to their seemingly exponential cost in $d$. This spurred the creation of indirect Monte Carlo (MC) algorithms for probabilistic turbulence simulations [8]. These schemes have proven highly successful, enabling advanced turbulent combustion simulations involving thousands of CPU cores [9, 10]. However, the randomness and slow convergence characterising MC methods can be avoided by directly solving the underlying Fokker-Planck equations.
@@
It is not just probabilistic turbulence calculations that are hindered by the curse of dimensionality: quantum many-body systems are described by states whose sizes also grow exponentially (in the number of particles). However, physically-relevant quantum states are known to be highly structured [11]. Such structure can be exploited to compress the states into approximate, but highly accurate, polynomially large representations known as tensor networks (TNs). TN algorithms allow efficiently evolving these states and analyzing their physical properties without ever leaving the compressed TN representation [12, 13, 14, 15], and have enabled the simulation of otherwise intractable quantum systems like superconductors, ferromagnets and quantum computers [16, 17, 18, 19, 20, 21, 22, 23, 24]. Recently, the TN formalism has begun spreading beyond quantum physics [25, 26, 27, 28, 29, 30].
@@
Decades of empirical experience indicates that $f$ is also highly structured. For instance, in homogeneous turbulence, velocities $\mathbf{U}$ are often distributed normally [2], whereas mass-fractions $\Phi_{\alpha}$ have been observed to follow normal, exponential and beta distributions in non-reactive flows [31]. In more complicated reacting flows, the PDFs generally cannot be so simply parameterised [32], although they remain smoother and more predictable than the underlying flow-fields [33].
@@
This work shows that turbulence PDFs indeed contain structure which is readily exploitable through TNs: using a simple TN known as the “matrix product state” (MPS) ansatz to encode $f$ in a highly compressed format allows us to formulate a scheme for cheaply and directly solving the governing Fokker-Planck equations. When the PDF structure is well-matched to the MPS ansatz, the time-evolution costs just $\sim d\log M$; compared to the $\sim M^{d}$ scaling of standard FD schemes. We demonstrate the advantage by looking at the following turbulent flow.
Figure 1: High-dimensional PDF of a flow undergoing turbulent mixing revealed by TN simulation. Here the Fokker-Planck Eq. (3) is solved for a PDF $f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ over chemical mass-fractions $\varphi_{1},\varphi_{2}$, at $C_{\Omega}=1$, $\text{Da}=0$ in the presence of a $\text{Pe}=10^{3}$ velocity field $\braket{\mathbf{U}}$ characterised by vortices and a jet along $x_{1}$ (Methods, Flow case definition). The 5-dimensional $f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ is represented by a MPS ansatz at $\chi=128$, on a $128^{\times 5}$ grid and is visualised here for $\mathbf{x}/l_{0}=(\frac{1}{2},1,\frac{1}{2})$ and $\mathbf{x}/l_{0}=(0,\frac{1}{2},1)$ at times $t/T_{0}=0,0.125,1,2$ in the left and right columns, while corresponding mean mass-fraction ratios $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ are shown in the centre.
@@ Results
Consider an incompressible, 3D turbulent flow in which two chemical species are irreversibly reacting: $A+B\rightarrow\text{Products}$. In this system, two chemical species (of mass-fractions $\Phi_{1},\Phi_{2}$) are stirred by a velocity field $\mathbf{U}=\mathbf{U}(\mathbf{x},t)$ in 3D space $\mathbf{x}$ across time $t$. For the sake of simplicity, we here consider only the statistics of the $\Phi_{1},\Phi_{2}$ scalar fields by assuming that the large-scale statistical features of the hydrodynamics are known a priori, while modelling the subgrid-scale (SGS), turbulent velocity fluctuations using large eddy simulation (LES), per current best practices [10]. Doing so eliminates the randomness of $\mathbf{U}$ and reduces the dimensionality to $d=5+1$. Now, $f=f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ describes the statistics of the mass-fraction fluctuations, which provide the mean mass-fractions $\braket{\Phi_{1}},\braket{\Phi_{2}}$ through
@@
Such PDFs are known as “filtered density functions” [3, 34]. Deriving the equation governing $f$ requires SGS closure modelling. Using popular closure models [34] gives the Fokker-Planck PDE:
@@
$\begin{split}&\frac{\partial f}{\partial t}+\braket{U_{i}}\frac{\partial f}{\partial x_{i}}-\frac{\partial}{\partial x_{i}}\left[(\gamma+\gamma_{\text{SGS}})\frac{\partial f}{\partial x_{i}}\right]=\\ &\frac{\partial}{\partial\varphi_{\alpha}}\left[\Omega_{\text{mix}}\left(\varphi_{\alpha}-\braket{\Phi_{\alpha}}\right)f\right]-\frac{\partial}{\partial\varphi_{\alpha}}(S_{\alpha}f).\end{split}$ (3)
@@
Here $\braket{\mathbf{U}}(\mathbf{x},t)$ is the large-scale (or, “filtered”) mean hydrodynamic field across $\mathbf{x}\in[0,l_{0})^{\times 3}$, $t\in[0,2T_{0})$, which is set to be a jet flow combined with a Taylor-Green vortex of amplitude $u_{0}=l_{0}/T_{0}$ (Methods, Flow case definition).
@@
The left hand side of Eq. (3) denotes the PDF-transport in space and time. The first term is the rate of temporal change, and the second term represents convection by the mean velocity field. The third represents the influence of the molecular ($\gamma$) and SGS diffusion [$\gamma_{\text{SGS}}(\mathbf{x},t)$] coefficients: the former sets the Peclet number $\text{Pe}=u_{0}l_{0}/\gamma$, and the latter is modelled via the Smagorinsky [35] closure $\gamma_{\text{SGS}}=C_{s}\frac{\Delta_{\ell}^{2}}{2}\sqrt{\sum_{ij}\left(\frac{\partial U_{i}}{\partial x_{j}}+\frac{\partial U_{j}}{\partial x_{i}}\right)^{2}}$, with $C_{s}$ an empirical constant and $\Delta_{\ell}$ the LES filter width (both are specified in Methods, Flow case definition).
@@
The right hand side of Eq. (3) designates transport in the composition space [“composition” since the $\varphi_{1},\varphi_{2}\in[0,1)^{\times 2}$ mass-fractions define the composition of the fluid]. The first term represents scalar mixing from the SGS turbulence, and is modelled via the popular least-mean-square estimation (LMSE) [36] closure $\Omega_{\text{mix}}=C_{\Omega}\frac{\gamma+\gamma_{\text{SGS}}}{\Delta_{\ell}^{2}}$, with $C_{\Omega}$ the SGS mixing-rate. The final term denotes the effects of chemical reaction. For the binary reaction scheme considered here, $S_{1}=S_{2}=-C_{r}\varphi_{1}\varphi_{2}$, where $C_{r}$ denotes the reaction rate that defines the Damköhler number $\text{Da}=C_{r}l_{0}/u_{0}$.
@@
To solve Eq. (3), we discretise $f$ at every point in time on a $M=128,d=5$ Cartesian grid, but parameterise it as an MPS-network using far fewer variables than the $128^{5}$ gridpoints resolving it. This allows us to use a simple Runge-Kutta 2, FD scheme (Methods, Finite difference discretisation) to solve Eq. (3) and time-evolve the MPS-PDF.
@@
An initial MPS simulation (Methods, Matrix product state algorithm) is performed in Fig. 1 of a purely mixing flow without chemical reactions ($\text{Da}=0$). The PDF is illustrated at two points in $\mathbf{x}$ along with the scalar-ratio $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ at four different times, showing how the initially orderly, unmixed flow state is driven towards a fully-mixed $\braket{\Phi_{1}}/\braket{\Phi_{2}}\approx 1$ state by SGS and large-scale convective and diffusive mixing. The SGS mixing leads to the PDF concentrating, whilst the diffusion and mean-flow convection induces multi-modality in the PDF. The MPS simulation is highly accurate (Fig. 2b), yet the number of variables parameterising the PDF (NVPP) is only $\mathcal{O}(1/10^{5})$ of an equivalent, classically-implemented FD scheme [Methods, Eq. (12)].
@@
In our MPS encoding, the discretised, high-dimensional $f(\varphi_{1},\varphi_{2},x_{1},x_{2},x_{3})$ is decomposed into a 1D chain of tensors, where the $\varphi_{1},\varphi_{2},x_{1},x_{2},x_{3}$ dimensions are sequentially mapped to tensors from left to right, with each dimension itself decomposed into multiple tensors lengthscale-by-lengthscale (analogously to the “sequential, serial” ordering in [27, Fig. 1] and [37, Eq. 9]). This encoding exploits two separate structures that characterise the solution of Eq. (3): firstly, the general smoothness of turbulence PDFs; secondly, that the different dimensions of $f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ are unlikely to be strongly coupled at low $C_{\Omega}$ (Suppl. Sec. 1).
@@
Figure 2: Accuracy convergence of TN algorithm. The influence of $\chi,C_{\Omega}$ and Da on the accuracy of the MPS simulation is outlined here. Parts a and b contrast numerically exact means against those extracted from the MPS algorithm. In a, the ratio $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ is visualised at $x_{3}/l_{0}=\frac{1}{2}$ for times $t/T_{0}=0.125,0.25,1,2$, top to bottom. The leftmost column corresponds to the exact solution [which can only be practically computed for $\text{Da}=0$ through Eq. (9)], while the next six columns come from MPS simulations at varying $\chi,C_{\Omega}$. The differences between the exact and MPS solutions are quantified in the lower b plot; the upper b plot shows how well the total species amounts $\overline{\braket{\Phi_{\alpha}}}$ [see Eq. (4)] are preserved through the simulation. In c, the RMSE in two basic statistics is computed: the difference in species consumption $\overline{\braket{\Phi_{1}}-\braket{\Phi_{2}}}$, which should always equal zero, and the space-averaged norm $\overline{\langle 1\rangle}$, which should always equal one. All RMSEs are mathematically defined in Methods, Error measures.
@@
Increasing $C_{\Omega}$ is expected to lead to higher coupling between the different dimensions of $f$, which reduces the efficiency of our MPS encoding; i.e., increasing $C_{\Omega}$ requires an increased $\chi$ to maintain accuracy. To verify, we first set $\text{Da}=0$ because this allows us to accurately compute $\braket{\Phi_{\alpha}}$ independently of Eq. (3) (see Methods, Moment equations) and benchmark the accuracy of the computed MPS-PDF across $C_{\Omega},\chi$. The benchmark is shown in Figs. 2a and 2b. The $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ ratios in Fig. 2a depict how the MPS-PDF means approach their numerically exact equivalent when $\chi$ increases and $C_{\Omega}$ decreases. All the cases, including the ones with lowest accuracy, correctly trend towards a fully-mixed equilibrium state where $\braket{\Phi_{1}}/\braket{\Phi_{2}}\approx 1$. Figure 2b quantitatively shows that the root-mean-square-error (RMSE) in terms of both the Reynolds-averaged mean mass-fraction
@@
The maximal bond-dimension $\chi$ not only sets the accuracy of the MPS simulation, but also determines the computational cost. Because our MPS algorithm (Methods, Matrix product state algorithm) implements a finite difference method within the MPS framework, it must perform the MPS-equivalent of operations like element-wise, matrix-matrix and matrix-vector multiplications, matrix and vector additions and subtractions, and inner and outer products, in addition to MPS-specific operations like singular values and QR decompositions to enforce the maximal bond-dimension and ensure the MPS stays in the numerically manageable “canonical form” [14, 15]. It is straightforward to show [38] these MPS operations all cost $O(\chi^{q}d\log M)$ asymptotically, with $q\in\mathbb{Z}_{>0}$ depending on the operation.
The element-wise multiplication operation is the most expensive at $q=4$ [38, Section 4.6], making the asymptotic complexity of our scheme as a whole $O(\chi^{4}d\log M)$ per timestep. Thus for $M_{t}$ timesteps, the total cost of the time-evolution will approach $O(M_{t}\chi^{4}d\log M)$ at very large $\chi$; although in practice for small and intermediate $\chi$, the empirical cost scales much milder (Suppl. Sec. 2). In comparison, standard FD schemes are exponentially more expensive in $d$, costing $O(M_{t}M^{d})$.
There is also the question of preparing initial states and extracting statistics. Regarding the former, the 3D $\braket{\mathbf{U}}(\mathbf{x}),\gamma_{\text{SGS}}(\mathbf{x})$ and $d$-dimensional $f(\varphi_{1},\varphi_{2};\mathbf{x},t=0)$ can be computed using either the prolongation method (see [38, Sec. 4.4] and [39]) or the tensor-cross algorithm [40, 41, 30, 42], both at $O(\chi^{3}d\log M)$ cost. As for the latter, computing expectation values boils down to doing the MPS-equivalent of matrix-vector multiplication and inner products, which are, as noted previously, inexpensive and straightforward operations. For instance, at any given timestep, the 3D mean $\braket{\Phi_{\alpha}}$ can be extracted from $f$ at $O(M^{3}\chi^{2}d\log M)$ complexity, while the cost is $O(\chi^{2}d\log M)$ for the scalar $\overline{\braket{\Phi_{\alpha}}}$.
@@
$\overline{\Upsilon_{12}}=\overline{\braket{\Phi_{1}\Phi_{2}}-\braket{\Phi_{1}}\braket{\Phi_{2}}}.$ (6)
At initial times, $-\overline{\Upsilon_{12}}$ increases due to large gradients in $\braket{\Phi_{\alpha}}$, followed by a decrease due to mixing (with higher mixing-rates leading to a faster decay). The ratio $R_{12}/(R_{12}+\overline{\Upsilon_{12}})$ is consistently above one half, implying most of the energy of the eddies is resolved during the simulations. The statistical trends observed in Fig. 4 are consistent with those reported in turbulence literature [34].
Figure 4: Statistics extracted from PDF for different flow parameters. In the first row, the total amount of the first species is plotted. The next row is of the Reynolds-averaged covariance $R_{12}$, while the final row quantifies the space-averaged covariance $\overline{\Upsilon_{12}}$. These quantities are defined in Eqs. 4, 5 and 6. The PDFs are computed at $\chi=128$.

### v2

@@ Abstract
Predicting the dynamics of turbulent fluids has been an elusive goal for centuries. Even with modern computers, anything beyond the simplest turbulent flows are too chaotic and multi-scaled to be directly simulatable. An alternative is to treat turbulence probabilistically, viewing flow properties as random variables distributed according to joint probability density functions (PDFs). Such PDFs are neither chaotic nor multi-scale, yet remain challenging to simulate due to their high dimensionality. Here we overcome the dimensionality problem by encoding turbulence PDFs as highly compressed “tensor networks” (TNs). This enables single CPU core simulations that would otherwise be impractical even with supercomputers: for a $5+1$ dimensional PDF of a chemically reactive turbulent flow, we achieve reductions in memory and computational costs by factors of $\mathcal{O}(10^{6})$ and $\mathcal{O}(10^{3})$, respectively, compared to standard finite-difference algorithms. A future path is opened towards something heretofore thought infeasible: directly simulating high-dimensional PDFs of both turbulent flows, and other chaotic systems that can usefully be described probabilistically.
@@ Introduction
Despite the simple and deterministic physical laws governing it, turbulence remains an inherently complex and chaotic phenomenon. It is characterised by large numbers of eddies interacting in intricate and nonlinear ways across wide ranges of spatial and temporal scales, leading to the emergence of chaos. Making matters worse, practically important turbulent flows (e.g. fuel-oxidiser mixtures in combustion) often involve multiple chemically-reacting species, which introduces additional nonlinearities and scales. The presence of chaos prohibits predicting the exact dynamics of turbulent flow-fields over long periods of time, while the multi-scaled nature of the flow-fields makes their simulation immensely expensive due to the need to solve sets of coupled partial differential equations (PDEs) on very fine grids.
However, for practical applications it is rarely necessary to know the precise state of a turbulent flow-field at every point in space-time. Rather, one is typically more interested in far slower-varying statistical quantities where the fluctuations are averaged out (such as the lift and drag of an aeroplane, or the rate of product formation in a chemical process). In the statistical description of turbulence, variables like velocities $\mathbf{U}$, chemical mass-fractions $\Phi_{\alpha}$, temperatures, etc., are treated as random variables (RVs) distributed according to some one-point, one-time joint probability density function (PDF)  (?)
@@
across space $\mathbf{x}$ and time $t$, with $\mathbf{u},\varphi_{1}$ being sample-space variables corresponding to $\mathbf{U},\Phi_{1}$. The trajectory of $f$ completely describes the one-point, one-time statistics of the flow dynamics  (?); which are the central quantities of interest in practical engineering calculations.
@@
The time-evolution of $f$ is modelled by Fokker-Planck PDEs that are straightforward to derive  (?, ?, ?), but hard to solve. If $f$ is $d$-dimensional, assigning $M$ points for each dimension results in a total of $M^{d}$ gridpoints. Given that $d$ can be as high as $\mathcal{O}(10^{3})$ in realistic flows  (?, ?), direct schemes like finite-differences (FD) or volumes were long ago dismissed as computationally infeasible  (?) due to their seemingly exponential cost in $d$. This spurred the creation of indirect Monte Carlo (MC) algorithms for probabilistic turbulence simulations  (?). These schemes have proven highly successful, enabling advanced turbulent combustion simulations involving thousands of CPU cores  (?, ?). However, the randomness and slow convergence characterising MC methods can be avoided by directly solving the underlying Fokker-Planck equations.
@@
It is not just probabilistic turbulence calculations that are hindered by the curse of dimensionality: quantum many-body systems are described by states whose sizes also grow exponentially (in the number of particles). However, physically-relevant quantum states are known to be highly structured  (?). Such structure can be exploited to compress the states into approximate, but highly accurate, polynomially large representations known as tensor networks (TNs). TN algorithms allow efficiently evolving these states and analyzing their physical properties without ever leaving the compressed TN representation  (?, ?, ?, ?), and have enabled the simulation of otherwise intractable quantum systems like superconductors, ferromagnets and quantum computers  (?, ?, ?, ?, ?, ?, ?, ?, ?). Recently, the TN formalism has begun spreading beyond quantum physics  (?, ?, ?, ?, ?, ?, ?).
@@
Decades of empirical experience indicates that $f$ is also highly structured. For instance, in homogeneous turbulence, velocities $\mathbf{U}$ are often distributed normally  (?), whereas mass-fractions $\Phi_{\alpha}$ have been observed to follow normal, exponential and beta distributions in non-reactive flows  (?). In more complicated reacting flows, the PDFs generally cannot be so simply parameterised  (?), although they remain smoother and more predictable than the underlying flow-fields  (?).
@@
This work shows that the structure contained in turbulence PDFs is readily exploitable through TNs: using a simple TN known as the “matrix product state” (MPS) ansatz to encode $f$ in a highly compressed format allows us to formulate a scheme for cheaply and directly solving the governing Fokker-Planck equations. When the PDF structure is well-matched to the MPS ansatz, the time-evolution costs just $\sim d\log M$; while standard FD schemes scale as $\sim M^{d}$. We demonstrate the advantage by looking at the following turbulent flow.
@@ Results
Figure 1: High-dimensional PDF of a flow undergoing turbulent mixing revealed by TN simulation. Here the Fokker-Planck Eq. (3) is solved for a PDF $f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ over chemical mass-fractions $\varphi_{1},\varphi_{2}$, at $C_{\Omega}=1$, $\text{Da}=0$ in the presence of a $\text{Pe}=10^{3}$ velocity field $\braket{\mathbf{U}}$ characterised by vortices and a jet along $x_{1}$ (Materials and Methods, Flow case definition). The 5-dimensional $f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ is represented by a MPS ansatz at $\chi=128$, on a $128^{\times 5}$ grid and is visualised here for $\mathbf{x}/l_{0}=(\frac{1}{2},1,\frac{1}{2})$ and $\mathbf{x}/l_{0}=(0,\frac{1}{2},1)$ at times $t/T_{0}=0,0.125,1,2$ in the left and right columns, while corresponding mean mass-fraction ratios $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ are shown in the centre.
Consider an incompressible, 3D turbulent flow in which two chemical species are irreversibly reacting: $A+B\rightarrow\text{Products}$. In this system, two chemical species (of mass-fractions $\Phi_{1},\Phi_{2}$) are stirred by a velocity field $\mathbf{U}=\mathbf{U}(\mathbf{x},t)$ in 3D space $\mathbf{x}$ across time $t$. For the sake of simplicity, we here consider only the statistics of the $\Phi_{1},\Phi_{2}$ scalar fields by assuming that the large-scale statistical features of the hydrodynamics are known a priori, while modelling the subgrid-scale (SGS), turbulent velocity fluctuations using large eddy simulation (LES), per current best practices  (?). Doing so eliminates the randomness of $\mathbf{U}$ and reduces the dimensionality to $d=5+1$. Now, $f=f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ describes the statistics of the mass-fraction fluctuations, which provide the mean mass-fractions $\braket{\Phi_{1}},\braket{\Phi_{2}}$ through
@@
Such PDFs are known as “filtered density functions”  (?, ?). Deriving the equation governing $f$ requires SGS closure modelling. Using popular closure models  (?) gives the Fokker-Planck PDE:
@@
$\frac{\partial f}{\partial t}+\braket{U_{i}}\frac{\partial f}{\partial x_{i}}-\frac{\partial}{\partial x_{i}}\left[(\gamma+\gamma_{\text{SGS}})\frac{\partial f}{\partial x_{i}}\right]=\frac{\partial}{\partial\varphi_{\alpha}}\left[\Omega_{\text{mix}}\left(\varphi_{\alpha}-\braket{\Phi_{\alpha}}\right)f\right]-\frac{\partial}{\partial\varphi_{\alpha}}(S_{\alpha}f).$ (3)
@@
Here $\braket{\mathbf{U}}(\mathbf{x},t)$ is the large-scale (or, “filtered”) mean hydrodynamic field across $\mathbf{x}\in[0,l_{0})^{\times 3}$, $t\in[0,2T_{0})$, which is set to be a jet flow combined with a Taylor-Green vortex of amplitude $u_{0}=l_{0}/T_{0}$ (Materials and Methods, Flow case definition).
@@
The left hand side of Eq. (3) denotes the PDF-transport in space and time. The first term is the rate of temporal change, and the second term represents convection by the mean velocity field. The third represents the influence of the molecular ($\gamma$) and SGS diffusion [$\gamma_{\text{SGS}}(\mathbf{x},t)$] coefficients: the former sets the Peclet number $\text{Pe}=u_{0}l_{0}/\gamma$, and the latter is modelled via the Smagorinsky  (?) closure $\gamma_{\text{SGS}}=C_{s}\frac{\Delta_{\ell}^{2}}{2}\sqrt{\sum_{ij}\left(\frac{\partial U_{i}}{\partial x_{j}}+\frac{\partial U_{j}}{\partial x_{i}}\right)^{2}}$, with $C_{s}$ an empirical constant and $\Delta_{\ell}$ the LES filter width (both are specified in Materials and Methods, Flow case definition).
@@
The right hand side of Eq. (3) designates transport in the composition space [“composition” since the $\varphi_{1},\varphi_{2}\in[0,1)^{\times 2}$ mass-fractions define the composition of the fluid]. The first term represents scalar mixing from the SGS turbulence, and is modelled via the popular least-mean-square estimation (LMSE)  (?) closure $\Omega_{\text{mix}}=C_{\Omega}\frac{\gamma+\gamma_{\text{SGS}}}{\Delta_{\ell}^{2}}$, with $C_{\Omega}$ the SGS mixing-rate. The final term denotes the effects of chemical reaction. For the binary reaction scheme considered here, $S_{1}=S_{2}=-C_{r}\varphi_{1}\varphi_{2}$, where $C_{r}$ denotes the reaction rate that defines the Damköhler number $\text{Da}=C_{r}l_{0}/u_{0}$.
@@
To solve Eq. (3), we discretise $f$ at every point in time on a $M=128,d=5$ Cartesian grid, but parameterise it as an MPS-network using far fewer variables than the $128^{5}$ gridpoints resolving it. This allows us to use a simple Runge-Kutta 2, FD scheme (Materials and Methods, Finite-difference discretisation) to solve Eq. (3) and time-evolve the MPS-PDF.
@@
An initial MPS simulation (Materials and Methods, Matrix product state algorithm) is performed in Fig. 1 of a purely mixing flow without chemical reactions ($\text{Da}=0$). The PDF is illustrated at two points in $\mathbf{x}$ along with the scalar-ratio $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ at four different times, showing how the initially orderly, unmixed flow state is driven towards a fully-mixed $\braket{\Phi_{1}}/\braket{\Phi_{2}}\approx 1$ state by SGS and large-scale convective and diffusive mixing. The SGS mixing leads to the PDF concentrating, whilst the diffusion and mean-flow convection induces multi-modality in the PDF. The MPS simulation is highly accurate (Fig. 2b), yet the number of variables parameterising the PDF (NVPP) is only $\mathcal{O}(1/10^{5})$ of an equivalent, classically-implemented FD scheme [Materials and Methods, Eq. (12)].
@@
In our MPS encoding, the discretised, high-dimensional $f(\varphi_{1},\varphi_{2},x_{1},x_{2},x_{3})$ is decomposed into a 1D chain of tensors, where the $\varphi_{1},\varphi_{2},x_{1},x_{2},x_{3}$ dimensions are sequentially mapped to tensors from left to right, with each dimension itself decomposed into multiple tensors lengthscale-by-lengthscale [analogously to the “sequential, serial” ordering in  (?, Fig. 1) and  (?, Eq. 9)]. This encoding exploits two separate structures that characterise the solution of Eq. (3): firstly, the general smoothness of turbulence PDFs; secondly, that the different dimensions of $f(\varphi_{1},\varphi_{2};\mathbf{x},t)$ are unlikely to be strongly coupled at low $C_{\Omega}$, because for $C_{\Omega}=0$ the PDF is separable (Supplementary Text, Separability of Fokker-Planck equation).
@@
Figure 2: Accuracy convergence of TN algorithm. The influence of $\chi,C_{\Omega}$ and Da on the accuracy of the MPS simulation is outlined here. Panels A and B contrast numerically exact means against those extracted from the MPS algorithm. In A, the ratio $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ is visualised at $x_{3}/l_{0}=\frac{1}{2}$ for times $t/T_{0}=0.125,0.25,1,2$, top to bottom. The leftmost column corresponds to the exact solution [which can only be practically computed for $\text{Da}=0$ through Eq. (9)], while the next six columns come from MPS simulations at varying $\chi,C_{\Omega}$. The differences between the exact and MPS solutions are quantified in the lower B plot; the upper B plot shows how well the total species amounts $\overline{\braket{\Phi_{\alpha}}}$ [see Eq. (4)] are preserved through the simulation. In C, the RMSE in two basic statistics is computed: the difference in species consumption $\overline{\braket{\Phi_{1}}-\braket{\Phi_{2}}}$, which should always equal zero, and the space-averaged norm $\overline{\langle 1\rangle}$, which should always equal one. All RMSEs are mathematically defined in Materials and Methods, Error measures.
@@
Increasing $C_{\Omega}$ is expected to lead to higher coupling between the different dimensions of $f$, which reduces the efficiency of our MPS encoding; i.e., increasing $C_{\Omega}$ requires an increased $\chi$ to maintain accuracy. To verify, we first set $\text{Da}=0$ because this allows us to accurately compute $\braket{\Phi_{\alpha}}$ independently of Eq. (3) (see Materials and Methods, Moment equations) and benchmark the accuracy of the computed MPS-PDF across $C_{\Omega},\chi$. The benchmark is shown in Figs. 2a and 2b. The $\braket{\Phi_{1}}/\braket{\Phi_{2}}$ ratios in Fig. 2a depict how the MPS-PDF means approach their numerically exact equivalent when $\chi$ increases and $C_{\Omega}$ decreases. All the cases, including the ones with lowest accuracy, correctly trend towards a fully-mixed equilibrium state where $\braket{\Phi_{1}}/\braket{\Phi_{2}}\approx 1$. Figure 2b quantitatively shows that the root-mean-square-error (RMSE) in terms of both the Reynolds-averaged mean mass-fraction
@@
The maximal bond-dimension $\chi$ not only sets the accuracy of the MPS simulation, but also determines the computational cost. Because our MPS algorithm (Materials and Methods, Matrix product state algorithm) implements a finite-difference method within the MPS framework, it must perform the MPS-equivalent of operations like element-wise, matrix-matrix and matrix-vector multiplications, matrix and vector additions and subtractions, and inner and outer products, in addition to MPS-specific operations like singular values and QR decompositions to enforce the maximal bond-dimension and ensure the MPS stays in the numerically manageable “canonical form”  (?, ?). It is straightforward to show  (?) these MPS operations all cost $O(\chi^{q}d\log M)$ asymptotically, with $q\in\mathbb{Z}_{>0}$ depending on the operation.
The element-wise multiplication operation is the most expensive at $q=4$  (?, Section 4.6), making the asymptotic complexity of our scheme as a whole $O(\chi^{4}d\log M)$ per timestep. Thus for $M_{t}$ timesteps, the total cost of the time-evolution will approach $O(M_{t}\chi^{4}d\log M)$ at very large $\chi$; although in practice for small and intermediate $\chi$, the empirical cost scales much milder (Supplementary Text, Empirical computational cost). In comparison, standard FD schemes are exponentially more expensive in $d$, costing $O(M_{t}M^{d})$.
There is also the question of preparing initial states and extracting statistics. Regarding the former, the 3D $\braket{\mathbf{U}}(\mathbf{x}),\gamma_{\text{SGS}}(\mathbf{x})$ and $d$-dimensional $f(\varphi_{1},\varphi_{2};\mathbf{x},t=0)$ can be computed using either the prolongation method [see  (?, Sec. 4.4) and  (?)] or the tensor-cross algorithm  (?, ?, ?, ?, ?), both at $O(\chi^{3}d\log M)$ cost. As for the latter, computing expectation values boils down to doing the MPS-equivalent of matrix-vector multiplication and inner products, which are, as noted previously, inexpensive and straightforward operations. For instance, at any given timestep, the 3D mean $\braket{\Phi_{\alpha}}$ can be extracted from $f$ at $O(M^{3}\chi^{2}d\log M)$ complexity, while the cost is $O(\chi^{2}d\log M)$ for the scalar $\overline{\braket{\Phi_{\alpha}}}$.
@@
$\overline{\text{Y}_{12}}=\overline{\braket{\Phi_{1}\Phi_{2}}-\braket{\Phi_{1}}\braket{\Phi_{2}}}.$ (6)
At initial times, $-\overline{\text{Y}_{12}}$ increases due to large gradients in $\braket{\Phi_{\alpha}}$, followed by a decrease due to mixing (with higher mixing-rates leading to a faster decay). The ratio $R_{12}/(R_{12}+\overline{\text{Y}_{12}})$ is consistently above one half, implying most of the energy of the eddies is resolved during the simulations. The statistical trends observed in Fig. 4 are consistent with those reported in turbulence literature  (?).
Figure 4: Statistics extracted from PDF for different flow parameters. In the first row, the total amount of the first species is plotted. The next row is of the Reynolds-averaged covariance $R_{12}$, while the final row quantifies the space-averaged covariance $\overline{\text{Y}_{12}}$. These quantities are defined in Eqs. 4, 5 and 6. The PDFs are computed at $\chi=128$.